OpenStudy (anonymous):

wat is -21/50 / 9/15 plez show work

OpenStudy (anonymous):

Did you mean to say\[-21/50 \div 9/15?\]

OpenStudy (anonymous):

To divide fractions you can multiply the reciprocal of one of the fractions times the other fraction, or in other words you can flip one fraction (it doesn't matter which one) and multiply it by the other (without flipping it). Like this: \[-21/50 \times 15/9.\] You would then need to multiply in the usual way (-21 times 15 and 50 times 9).

OpenStudy (anonymous):

By the way the answer would be -315/450 or -7/10 or -.7

OpenStudy (anonymous):

ITS DIVISION -21/50 DIVIDED BY 9/15

OpenStudy (anonymous):

Yeah, that is the same as -21/50 times 15/9. Just as (2/1) which is 2 by the way divided by (2) so all you do is multiply the flipped fraction (2/1)(1/2)=(2*1)/(1*2)=2/2=1